OpenStudy (nirmalnema):

what is the sq rt of i (i = sq rt of -1)

OpenStudy (zzr0ck3r):

(a + bi)^2 = i

OpenStudy (mimi_x3):

\[\sqrt{i} = (i)^{1/2} \] \[z = r(cos\theta + isin\theta)\] \[i => z = 1(cos(\frac{pi}{2})+ isin\frac{pi}{2})\]  using demoivre's \[\sqrt{i} => (cos(pi/2) + isin(pi/2))^{1/2}\] \[z=(cos(\frac{pi}{4})+isin(\frac{pi}{4})\]